### v1

# Continuous-time Data-driven Barrier Certificate Synthesis

Abstract
We consider the problem of verifying safety for continuous-time dynamical systems. Developing upon recent advancements in data-driven verification, we use only a finite number of sampled trajectories to learn a barrier certificate, namely a function which verifies safety. We train a safety-informed neural network to act as this certificate, with an appropriately designed loss function to encompass the safety conditions. In addition, we provide probabilistic generalisation guarantees from discrete samples of continuous trajectories, to unseen continuous ones. Numerical investigations demonstrate the efficacy of our approach and contrast it with related results in the literature. For the purpose of Open Access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
1 Introduction
Ensuring the safety of Continuous-time dynamical systems is of critical importance in an increasingly autonomous world [8, 15, 19]. As it is often infeasible to model system behaviour precisely, and making direct use of system data to verify behaviour is of interest [1, 12].
A technique to verify properties of dynamical systems involves discretising the state space [3], under approximation guarantees, and verifying the resulting model. Alternatively, the use of certificates [2, 18] allows one to analyse directly the continuous-state system. These certificates map system’s states to real values, and exhibit certain properties that are relevant for analysis: here in particular we construct safety certificates for continuous-time systems, but extensions to more complex certificates are possible  [19].
There are a number of techniques for synthesising such certificates. In the case that an exact model is known, one can use a polynomial function as a certificate to formulate a convex sum-of-squares problem [17]. Recent work in this area investigated the use of neural networks as certificates [8], which represent a class of general function approximators.
When obtaining an exact model is infeasible, we turn to data-driven techniques. One method for employing data in certificate synthesis is through the use of state pairs (i.e. states, and next-states), sampled from across the domain of interest. Such techniques are investigated in [15] for deterministic systems, and in [21] for stochastic systems. Both these works make use of the techniques in [11, 14] to bound the distance between what is referred to as a robust program, and its sample based counterpart. As discussed in [14, Remark 3.9], such techniques exhibit an exponential growth in the dimension of the sampling space (here the state space), and also require access to the underlying probability distribution.
Alternatively, one can consider using entire trajectories as samples, hence only requiring access to realistic runs of the system. This technique is explored in [19] for discrete-time systems. This work leverages the so-called scenario approach [5, 10] and the notion of compression [13, 7], and provides a constructive algorithm for the general pick-to-learn framework [16], to provide a probably approximately correct (PAC) bound on the correctness of certificates for newly sampled trajectories. These techniques are restricted to discrete-time systems, since they perform calculations on entire trajectories, which is infeasible in continuous-time. Hence, here we develop techniques for continuous-time systems by analysing samples from complete trajectories and bounding the difference between the safety of these approximations and the complete trajectories.
Our main contribution is the synthesis of neural barrier certificates for continuous-time dynamical systems, accompanied by PAC generalisation guarantees. These certificates are useful per se as a proxy for safety, and open the road for their use in control synthesis, an area of ongoing research.
Notation. We use $\{\xi_{k}\}_{k=0}^{K}$ to denote a sequence indexed by $k\in\{0,1,\dots,K\}$. $B\models\psi$ defines condition satisfaction i.e., it evaluates to true if the quantity $B$ on the left satisfies the condition $\psi$ on the right, e.g., $x=1\models x>0$ evaluates to true and $x=-1\models x>0$ evaluates to false. Using ⊧$\not\models$̸ represents the logical inverse of this (i.e., condition dissatisfaction). By $(\forall\xi\in\Xi)B\models\psi(\xi)$ we mean that some quantity $B$ satisfies a condition $\psi$ which, in turn, depends on some parameter $\xi$, for all $\xi\in\Xi$. We use $\xi_{[0,k]}$ to refer to a (possibly infinite) subsequence $\{\xi_{0},\dots,\xi_{k}\}$ of a sequence.
2 Certificates
In this work, we focus on barrier certificates, but emphasise that our techniques naturally extend to the more complex certificates in [19]. To this end, we begin by defining a dynamical system, before considering the barrier certificate and associated safety property it verifies.
2.1 Continuous-Time Dynamical Systems
We consider a bounded state space $X\subseteq\mathbb{R}^{n}$, and a dynamical system whose evolution starts at an initial state $x(0)\in X_{I}$, where $X_{I}\subseteq X$ denotes the set of possible initial conditions. From an initial state, we uncover a finite trajectory, i.e., a continuous sequence of states $\xi=\{x(t)\}_{t\in[0,T]}$, where $T\in\mathbb{R}$ and $x(t)\colon t\rightarrow\mathbb{R}^{n}$, by following the dynamics
$\dot{x}(t)=f(x(t)).$ (1)
We only require $f\colon X\rightarrow\mathbb{R}^{n}$ to be Lipschitz continuous, thus enforcing unique solutions. The set of all possible trajectories $\Xi\subseteq X_{I}\times X^{(0,T]}$ is then the set of all trajectories starting from the initial set $X_{I}$.
In Section 3, we discuss how to use a finite set of trajectories to build safety certificates, and accompany them with generalization guarantees with respect to their validity for any trajectory. To allow for the synthesis of such certificate, as discussed in Section 4, we must store and perform calculations on such trajectories that play the role of samples. For this reason, we consider a finite $T$, and take samples along the continuous trajectory. Hence, we define the time-discretised trajectory
$\tilde{\xi}=\{x(t)\}_{t\in\{0,t_{1},\dots,t_{M}\}}\in\tilde{\Xi}\subseteq X_{I}\times X^{M},$ (2)
for $M$ sampled time steps $t_{1},\dots,t_{M}$.
2.2 Safety Property
We use $\phi(\xi)$ to refer to the safety property under study, evaluated on a trajectory $\xi\in\Xi$, defined as follows.
Property 1 (Safety)
Consider (1), and let $X_{I},X_{U}\subset X$ with $X_{I}\cap X_{U}=\varnothing$ denote an initial and an unsafe set, respectively. If, for all $\xi\in\Xi$,
$\phi(\xi)\vcentcolon=\forall t\in[0,T],x(t)\notin X_{U},$
holds, then we say that $\phi$ encodes a safety property. $\Xi$ denotes the set of trajectories consistent with (1) and with initial state contained within $X_{I}$.
By the definition of $\phi$, it follows that verifying a system exhibits the safety property is equivalent to verifying all trajectories emanating from the initial set avoid the unsafe set for all time instances, until horizon $T$.
2.3 Barrier Certificate
We now define the relevant criteria necessary for a certificate $B$ to verify a safety property. We fix a time horizon $T<\infty$ and assume that $B$ is continuous, so that when considering the supremum/infimum of $B$ over $X$ (already assumed to be bounded) or over some of its subsets, this is well-defined. Consider:
$B(x)\leq 0,\forall x\in X_{I},$ (3)
$B(x)>0,\forall x\in X_{U},$ (4)
$\evaluated{\frac{dB}{dt}}_{x\in\xi}<\frac{1}{T}\Big{(}\inf_{x\in X_{U}}B(x)-\sup_{x\in X_{I}}B(x)\Big{)},$ (5)
where we may expand the Lie derivative as
$\frac{dB}{dt}=\frac{\partial B}{\partial x}\frac{dx}{dt}=\frac{\partial B}{\partial x}f(x),$ (6)
and hence recognise that this depends on the system dynamics $f(x)$. We place the following assumption on this Lie derivative, necessary for our analysis later.
Assumption 1 (Lie Derivative is Lipschitz)
Assume that $\frac{dB}{dt}$ is Lipschitz continuous.
Notice that even if $\inf_{x\in X_{U}}B(x)-\sup_{x\in X_{I}}B(x)>0$, i.e., in the case where the last condition encodes an increase of $B$ along the system trajectories, the system still avoids entering the unsafe set. This is established in the proof of Proposition 8 below.
A graphical representation of these conditions is found in Figure 3. The level set (with dashed line) is the set obtained when the certificate value is 00. The decrease condition then means that we never leave the sublevel set.
Denote by $\psi^{s}$ the conjunction of (3) and (4), and by $\psi^{\Delta}(\xi)$ the property in (5). Notice that the latter depends on $\xi$ as it relates to the derivative along a trajectory. We define a barrier certificate as follows.
Definition 1 (Property Verification & Certificates)
Given a safety property $\phi(\xi)$, and a function $B\colon\mathbb{R}^{n}\rightarrow\mathbb{R}$, let $\psi^{s}$ and $\psi^{\Delta}(\xi)$ be conditions such that, if
$\exists B\colon B\models\psi^{s}\wedge(\forall\xi\in\Xi)B\models\psi^{\Delta}(\xi)\implies\phi(\xi),\forall\xi\in\Xi,$
then the property $\phi$ is verified for all $\xi\in\Xi$. We then say that such a function $B$ is a barrier certificate.
In words, the implication of Definition 1 is that if a barrier certificate $B$ satisfies the conditions in $\psi^{s}$, as well as the conditions in $\psi^{\Delta}(\xi)$, for all $\xi\in\Xi$, then the safety property $\phi(\xi)$ is satisfied for all trajectories $\xi\in\Xi$.
Proposition 1 (Safety/Barrier Certificate)
A function $B\colon\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a safety/barrier certificate if
$B\models\psi^{s}\wedge(\forall\xi\in\Xi)B\models\psi^{\Delta}(\xi).$ (7)
Proof 2.1.
It suffices to show that satisfaction of (5) implies safety. Integrating (5) up to $t\leq T$, we obtain
$B(x(t))<B(x(0))+\frac{t}{T}\Big{(}\inf_{x\in X_{U}}B(x)-\sup_{x\in X_{I}}B(x)\Big{)}$
$\leq\frac{T-t}{T}\sup_{x\in X_{I}}B(x)+\frac{t}{T}\inf_{x\in X_{U}}B(x)$
$\leq\frac{t}{T}\inf_{x\in X_{U}}B(x)\leq\inf_{x\in X_{U}}B(x).$ (8)
where the second inequality is since $B(x(0))\leq\sup_{x\in X_{I}}B(x)$, as $x(0)\in X_{I}$. The third inequality is since $\sup_{x\in X_{I}}B(x)\leq 0$ due to (3), and the last one is since $t\leq T$. We thus have
$B(x(t))$ $<\inf_{x\in X_{U}}B(x),~{}t\in[0,T],$ (9)
i.e. the maximum value along a trajectory is less than the infimum over the unsafe region and hence $x(t)\notin X_{U},t=[0,T]$ (notice that $x(0)\notin X_{U}$ holds since $X_{I}\cap X_{U}=\varnothing$). The latter implies that all trajectories that start in $X_{I}$ avoid entering the unsafe set $X_{U}$, thus concluding the proof.
To synthesise one of these deterministic certificates, we require complete knowledge of the behaviour $f$ of the dynamical system, to allow us to evaluate the Lie derivative $\frac{dB}{dt}$. This may be impractical, and we therefore use data-driven techniques to learn a certificate.
3 Data-Driven Certificates
For our analysis, we will treat the initial state as random, distributed according to $\mathbb{P}$ (an appropriate probability space is defined; we gloss the technical details here in the interest of space). The support of $\mathbb{P}$ will be the set of admissible initial states (i.e. the initial set $X_{I}$).
To obtain our sample set, we consider $N$ independent and identically distributed (i.i.d.) initial conditions, sampled according to probability distribution $\mathbb{P}$, namely $\{x^{i}(0)\}_{i=1}^{N}\sim\mathbb{P}^{N}$. Initializing the dynamics from each of these initial states, we unravel a set of continuous-time trajectories $\{\xi^{i}\}_{i=1}^{N}$. Since there is no stochasticity in the dynamics, we can equivalently say that trajectories (generated from the random initial conditions) are distributed according to the same probabilistic law; hence, with a slight abuse of notation, we write $\xi\sim\mathbb{P}$. We impose the following assumption.
Assumption 1 (Non-concentrated Mass).
Assume that $\mathbb{P}\{\xi\}=0$, for any $\xi\in\Xi$.
3.1 Problem Statement
Since we are now dealing with a sample-based problem, we will construct probabilistic certificates and hence probabilistic guarantees on the satisfaction of a given property.
Denote by $B_{N}$ a barrier certificate, we introduce the subscript $N$ to emphasize that this certificate is constructed on the basis of time-discretised sampled trajectories $\{\tilde{\xi}^{i}\}_{i=1}^{N}$.
Problem 2 (Probabilistic Property Guarantee).
Consider $N$ sampled trajectories, and fix a confidence level $\beta\in(0,1)$. We seek a property violation level $\epsilon\in(0,1)$ such that
$\mathbb{P}^{N}$$\big{\{}\{\tilde{\xi}^{i}\}_{i=1}^{N}\in\tilde{\Xi}^{N}:~{}$
$\mathbb{P}\{\xi\in\Xi\colon B_{N}\not\models\psi^{s}\wedge\psi^{\Delta}(\xi)\}\leq\varepsilon\big{\}}\geq 1-\beta.$ (10)
In words, finding a solution to Problem 2 requires determining an $\epsilon\in(0,1)$, such that with confidence at least $1-\beta$, the probability that $B_{N}$ does not satisfy the condition $\psi^{s}\wedge\psi^{\Delta}(\xi)$ for another sampled continuous-time trajectory $\xi\in\Xi$ is at most equal to that $\epsilon$. As such, with a certain confidence, a certificate $B_{N}$ trained on the basis of $N$ sampled trajectories, will remain a valid certificate with probability at least $1-\epsilon$. Note that the outer probability in (10) (as well as in similar statements below) refers to the selection of discretized trajectories, $\{\tilde{\xi}^{i}\}_{i=1}^{N}$ (as these are used for training), while the inner probability refers to the selection of a continuous time one $\xi$, as this captures the desired generalization properties.
3.2 Probabilistic Guarantees
Consider a mapping $\mathcal{A}$ such that $B_{N}=\mathcal{A}(\{\xi^{i}\}_{i=1}^{N})$ as an algorithm that, based on $N$ samples, returns a certificate $B_{N}$. We call as compression set of such an algorithm any subset of the input that returns the same certificate. That is, a sample-subset $\mathcal{C}_{N}\subseteq\{\xi^{i}\}_{i=1}^{N}$ is a compression set if $\mathcal{A}(\mathcal{C}_{N})=\mathcal{A}(\{\xi^{i}\}_{i=1}^{N})$. In Algorithm 1, we provide a specific synthesis procedure through which $\mathcal{A}$ (and hence the certificate $B_{N}$) can be constructed. This algorithm is adapted from [19], with appropriate modifications to allow for continuous-time dynamics; we discuss this in the next section. Using the results of [19], and if we have access to samples of the time-derivative $f(x)$, we can provide a guarantee over the time-discretised trajectories, as stated next.
Theorem 3 (Probabilistic Guarantees [19]).
Consider Assumption 1, and let $B_{N}$ and $\mathcal{C}_{N}$ be the certificate and compression set, respectively, returned by Algorithm 1. Fix $\beta\in(0,1)$, and for $k<N$, let let $\varepsilon(k,\beta,N)$ be the (unique) solution to the polynomial equation in the interval $[k/N,1]$
$\frac{\beta}{2N}\sum_{m=k}^{N-1}$ $\frac{\binom{m}{k}}{\binom{N}{k}}(1-\varepsilon)^{m-N}$
$+\frac{\beta}{6N}\sum_{m=N+1}^{4N}\frac{\binom{m}{k}}{\binom{N}{k}}(1-\varepsilon)^{m-N}=1,$ (11)
while for $k=N$ let $\varepsilon(N,\beta,N)=1$. We then have that
$\mathbb{P}^{N}\big{\{}\{\tilde{\xi}^{i}\}_{i=1}^{N}\in\tilde{\Xi}^{N}:~{}$ (12)
$\mathbb{P}\{\tilde{\xi}\in\tilde{\Xi}\colon B_{N}\not\models\psi^{s}\wedge\psi^{\Delta}(\tilde{\xi}))\}\leq\varepsilon(C_{N},\beta,N)\big{\}}\geq 1-\beta,$
where $C_{N}=|\mathcal{C}_{N}|$ is the cardinality of the compression set.
Unfortunately, this does not provide us with a guarantee on the property satisfaction for the continuous-time trajectories, i.e. our trajectory may violate the safety property between sampled states $x(t_{i})$ and $x(t_{i+1})$, as the inner probability refers to a choice of $\tilde{\xi}$ as opposed to $\xi$.
If the Lipschitz continuity in Assumption 1 holds, we can additionally offer guarantees on newly sampled continuous-time trajectories, based only on time-discretised trajectories, and approximate derivatives. In this case, we require a tightening of the derivative condition in (5) by some value $d\in\mathbb{R}$, to ensure that the decrease condition is maintained between sample times. Denote this condition (over discretized trajectories) as $\psi^{\Delta}_{d}(\tilde{\xi})$, defined by the inequality
$\max_{k=1,\dots,M}$ $\frac{B(x(t_{k}))-B(x(t_{k-1}))}{t_{k}-t_{k-1}}$ (13)
$<\frac{1}{T}\Big{(}\inf_{x\in X_{U}}B(x)-\sup_{x\in X_{I}}B(x)\Big{)}-d.$
Define by $L_{B}$ and $L_{f}$ the Lipschitz constants of the certificate derivative $\frac{\partial B_{\theta^{\star}}}{\partial x}$ and of the dynamics $f(x)$ respectively, and by $\mathcal{M}_{B},\mathcal{M}_{f}$ bounds on their norms, namely $\sup_{x}\|\frac{\partial B_{\theta^{\star}}}{\partial x}\|\leq\mathcal{M}_{B}$ and $\sup_{x}\|f(x)\|\leq\mathcal{M}_{f}$. Then, the value $d$ is defined as
$d=\overline{t}\mathcal{M}_{f}\left(\mathcal{M}_{B}\mathcal{L}_{f}+\mathcal{M}_{f}\mathcal{L}_{B}\right),$ (14)
where $\overline{t}=\max_{k=1,\dots,M}(t_{k}-t_{k-1})$.
Theorem 4 (Continuous-Time Guarantees).
Consider the conditions of Theorem 3, Assumption 1 and (14). Then,
$\mathbb{P}^{N}\big{\{}\{\tilde{\xi}^{i}\}_{i=1}^{N}\in\tilde{\Xi}^{N}\colon$ (15)
$\mathbb{P}\{\xi\in\Xi\colon B_{N}\not\models\psi^{s}\wedge\psi^{\Delta}(\xi))\}\leq\varepsilon(C_{N},\beta,N)\big{\}}\geq 1-\beta.$
The proof of this is achieved by bounding the difference between the safety of the continuous-time trajectories and their time-discretised approximations, and is in the Appendix.
4 Certificate Synthesis
In order to learn a barrier certificate from samples, we consider a neural network, a well-studied class of function approximators that generalize well to a given task. Denote all tunable neural network parameters by a vector $\theta$. We then have that our certificate $B_{N}$ depends on $\theta$. For the results of this section, we simply write $B_{\theta}$ and drop the dependency on $N$ to ease notation.
Algorithm 1 Certificate Synthesis and Compression Set Computation
1:function $\mathcal{A}(\theta,\mathcal{D})$
2:     Set $k\leftarrow 1$triangleright Initialise iteration index
3:     Set $\mathcal{C}\leftarrow\varnothing$triangleright Initialise compression set
4:     Fix $L_{1}<L_{0}$ with $|L_{1}-L_{0}|>\eta$triangleright $\eta$ is any fixed tolerance
5:     while $l^{s}(\theta)>0$ do $\triangleright$ While sample-independent state loss is non-zero
6:         $g\leftarrow\nabla_{\theta}l^{s}(\theta)$triangleright Gradient of loss function
7:         $\theta\leftarrow\theta-\alpha g$triangleright Step in the direction of sample-independent gradient
8:
9:     while $|L_{k}-L_{k-1}|>\eta$ do $\triangleright$ Iterate until tolerance is met
10:         $\mathcal{M}\leftarrow\{\tilde{\xi}\in\mathcal{D}\colon L(\theta,\tilde{\xi})\geq\max_{\tilde{\xi}\in\mathcal{C}}L(\theta,\tilde{\xi})\}$triangleright Find maximal samples with loss greater than compression set loss
11:         $\overline{g}_{\mathcal{M}}\leftarrow\{\nabla_{\theta}L(\theta,\tilde{\xi})\}_{\tilde{\xi}\in\mathcal{M}}$triangleright Subgradients of loss function for $\tilde{\xi}\in\mathcal{M}$
12:         $\overline{\xi}_{\mathcal{C}}\in\operatorname*{arg\,max}_{\tilde{\xi}\in\mathcal{C}}L(\theta,\tilde{\xi})$triangleright Find a sample with maximum loss from $\mathcal{C}$
13:         $\overline{g}_{\mathcal{C}}\leftarrow\nabla_{\theta}L(\theta,\overline{\xi}_{\mathcal{C}})$triangleright Approximate subgradient of loss function for $\tilde{\xi}=\overline{\xi}_{\mathcal{C}}$
14:
15:         if $\exists\overline{g}\in\overline{g}_{\mathcal{M}}\colon\langle\overline{g},\overline{g}_{\mathcal{C}}\rangle\leq 0\wedge\overline{g}\neq 0$  then $\triangleright$ If there is a misaligned subgradient (take the maximum if multiple)
16:              $\theta\leftarrow\theta-\alpha\overline{g}$triangleright Step in the direction of misaligned subgradient
17:              $\mathcal{C}\leftarrow\mathcal{C}\cup\{\overline{\xi}\}$triangleright Update compression set with sample corresponding to $\overline{g}$
18:         else
19:              $\theta\leftarrow\theta-\alpha\overline{g}_{\mathcal{C}}$triangleright Step in the direction of approximate subgradient
20:
21:         $L_{k}\leftarrow\min\left\{L_{k-1},\max_{\tilde{\xi}\in\mathcal{D}}L(\theta,\tilde{\xi})\right\}$triangleright Update “running” loss value
22:         $k\leftarrow k+1$triangleright Update iteration index
23:     return $\theta,\mathcal{C}$
Figure 1: Graphical Representation of Algorithm 1.
4.1 Certificate and Compression Set Computation
We seek to minimize a loss function that encodes the barrier certificate conditions, with respect to the neural network parameters. To this end, for a $\tilde{\xi}\in\tilde{\Xi}$ and parameter vector $\theta$, let
$L(\theta,\tilde{\xi})=l^{\Delta}(\theta,\tilde{\xi})+l^{s}(\theta),$ (16)
represent an associated loss function comprised of sample-dependent loss $l^{\Delta}$, and sample-independent loss $l^{s}$
$l^{s}(V_{\theta})$ :$\vcentcolon=\frac{1}{|\mathcal{X}_{I}|}\sum_{x\in\mathcal{X}_{I}}\max\{0,V_{\theta}(x)\}$
$+\frac{1}{|\mathcal{X}_{U}|}\sum_{x\in\mathcal{X}_{U}}\max\{0,-V_{\theta}(x)\}.$
$l^{\Delta}(\theta,\tilde{\xi})$ :$\vcentcolon=-\frac{1}{T}\Big{(}\inf_{x\in\mathcal{X}_{U}}B(x)-\sup_{x\in\mathcal{X}_{I}}B(x)\Big{)}$
$+\max_{k=1,\dots,M}\frac{B(x(t_{k}))-B(x(t_{k-1}))}{t_{k}-t_{k-1}}.$
To instantiate these functions we consider a discrete set of grid-points on each sub-domain: $\mathcal{X}_{\overline{I}}$ is the set of points in the domain but outside the goal region, and $\mathcal{X}_{U}$ is the set of points in the unsafe set. These points are generated densely enough across the domain of interest, and hence offer an accurate approximation. Since these samples do not require access to the dynamics, we consider them separate to the sample-set $\{\tilde{\xi}^{i}\}_{i=1}^{N}$.
Consider the first summation in the sample-independent loss, if $B(x)<0$ then $\max\{0,B_{\theta}(x)\}=0$, i.e., no loss is incurred, implying satisfaction of (3). Under a similar reasoning, the other integral accounts for (4), respectively.
We impose the next mild assumption, a sufficient condition for termination of our algorithm.
Assumption 5 (Minimizers’ Existence).
For any $\{\tilde{\xi}\}_{i=1}^{N}$, and any non-empty $\mathcal{D}\subseteq\{\tilde{\xi}\}_{i=1}^{N}$, the set of minimizers of $\max_{\tilde{\xi}\in\mathcal{D}}L(\theta,\tilde{\xi}),$ is non-empty.
Note that, for a sufficiently expressive neural network, we can find a certificate $B$ which satisfies the state constraints and hence has a sample-independent loss of zero.
Algorithm 1 provides an inexact subgradient methodology to minimize the loss function, and to iteratively construct a compression set $\mathcal{C}$ (initially empty; see step $3$). We explain the main steps of this algorithm with reference to Figure 1. After an arbitrary parameter initialization, we follow the subgradient associated with the current sample in $\mathcal{C}$ (“blue” dot labeled by $3$). At this point, this step becomes inexact, as there would exist another sample resulting in a higher loss (“green” point labeled by $2$). Such a sample is in $\mathcal{M}$, step $9$ of Algorithm 1. However, the algorithm does not “jump” to the green point, as the condition in step $13$ of the algorithm is not yet satisfied. As such the algorithm performs inexact subgradient descent steps up to point $5$; this is the first instance where the condition in step $13$ is satisfied (there exists another constraint with opposite slope) and hence the algorithm “jumps” to a point with higher loss and subgradient of opposite sign. The “jumps” serve as an exploration step to investigate the non-convex landscape, while their number corresponds to the cardinality of the returned compression set. Such a procedure can be thought of as a constructive procedure for the general framework presented recently in [16] to construct compression sets.
In some cases, the parameter returned by Algorithm 1 may result in a value of the loss function greater than $-d$, and hence mean we are unable to verify safety. To achieve a lower loss, we make use of a sample-and-discarding procedure [6, 20], and introduce Algorithm 2 as an outer loop around Algorithm 1. This procedure leads to a lower loss, however the samples that are discarded have to be added to the compression set. Theorems 3 and 4 then apply, but with the compression set returned by Algorithm 2 instead.
To terminate our algorithm we require knowledge of $d$, but cannot calculate $d$ until we find $\theta^{\star}$. To resolve this, we propose two different approaches.
  1. 1.
    At every iteration $j$ calculate $d_{j}$ using the current best parameters $\theta_{j}$, terminate when $\max_{i}\left[L(\theta_{j},\tilde{\xi}^{i})\right]<d_{j}$.
  2. 2.
    Choose a parameter set $\Theta$, and take the supremum across the set to find an upper bound on $\mathcal{L}_{B},\mathcal{M}_{B}$, use these to calculate $d$.
To determine Lipschitz constants for neural networks we refer the reader to [4, 9, 22].
Algorithm 2 Compression Set Update with Discarding
1:Fix $\{\tilde{\xi}^{i}\}_{i=1}^{N}$
2:Set $\widetilde{\mathcal{C}}\leftarrow\varnothing$$\triangleright$ Initialise compression set
3:Set $\mathcal{D}\leftarrow\{\tilde{\xi}^{i}\}_{i=1}^{N}$triangleright Initialise “running” samples
4:while $(\max_{\tilde{\xi}\in\mathcal{D}}l^{\Delta}(\theta,\tilde{\xi})>-d)\bigvee(l^{s}(\theta)>0)$ do
5:     $\theta,\mathcal{C}\leftarrow$ $\mathcal{A}(\theta,\mathcal{D})$triangleright Call Algorithm 1
6:     $\widetilde{\mathcal{C}}\leftarrow\widetilde{\mathcal{C}}\cup\mathcal{C}$triangleright Update $\widetilde{\mathcal{C}}$
7:     $\mathcal{D}\leftarrow\mathcal{D}\setminus\widetilde{\mathcal{C}}$triangleright Discard $\widetilde{\mathcal{C}}$ from $\mathcal{D}$
8:return $\theta$, $\widetilde{\mathcal{C}}$
5 Numerical Results
We consider constructing a safety certificate for the nonlinear, two-dimensional jet engine model as considered in [15],
$\dot{x}_{1}(t)=-x_{2}(t)-\frac{3}{2}x_{1}^{2}(t)-\frac{1}{2}x_{1}^{3}(t),\;\;\dot{x}_{2}(t)=x_{1}(t).$ (17)
In Figure 3 we provide a graphical representation of the dynamics, sub-domains under study, the 00-level set produced by our certificate, and the level sets calculated by the methods in [15] (one lower bounding the unsafe set, the other upper bounding the initial set). We used $5$ independent repetitions (each with different multi-samples) of $1,000$ sampled trajectories and $367$ seconds of computation time (standard deviation $139$s), to obtain $\varepsilon=0.01492$ (standard deviation $0.00140$) with confidence to $0.99$. The methodology of [15] required $257149$ state pair samples and $5123$ seconds (standard deviation $449$s) of computation time to compute a barrier certificate with the same confidence (however, this holds deterministically). We estimate the Lipschitz constants using the methods in [23]. Figure 3 contains a 3D plot of the certificate.
Figure 2: Phase plane plot, initial and unsafe set for (17). The zero-level set for our certificate is dashed; level sets that bound the initial and unsafe sets (i.e. $\gamma$- and $\lambda$- level sets) in [15] are dotted.
Figure 3: Surface plot of the safety/barrier certificate, generated by our techniques, for the system of Figure 3.
Beyond these numerical results, we briefly discuss the theoretical differences between our approaches. The results in [15] offer a guarantee that, with a certain confidence, the safety property is always satisfied, in contrast to Theorem 4 where we provide such guarantees in probability (up to a quantifiable risk level $\varepsilon$). However, these “always” guarantees, albeit very useful, come with some challenges. Firstly, they are not applicable when part of the initial set is unsafe. Secondly, they implicitly require some knowledge of the underlying probability distribution. Finally, they are bound to an exponential growth with the dimension of the state space.
Related to this last point, we performed a comparison on the following four-dimensional system taken from [8].
$\dot{x}_{1}(t)$ $=x_{1}(t)+\frac{x_{1}(t)x_{2}(t)}{5}-\frac{x_{3}(t)x_{4}(t)}{2},$ (18)
$\dot{x}_{2}(t)$ $=\cos(x_{4}(t)),$
$\dot{x}_{3}(t)$ $=0.01\sqrt{|x_{1}(t)|},$
$\dot{x}_{4}(t)$ $=-x_{1}(t)-x_{2}(t)^{2}+\sin(x_{4}(t)).$
We calculate that the approach of [15] requires at least $10^{19}$ samples to return a confidence lower bounded by at most $10^{-30}$, which is not practically useful. In contrast, our techniques with only $100$ samples obtain a risk level of $\varepsilon=0.21450$ (standard deviation $0.00910$), confidence $1-10^{-5}$.
6 Conclusion
We have proposed a method for synthesis of neural-network certificates for continuous-time dynamical systems, based only on a finite number of trajectories from a system. Our numerical experiments demonstrate the efficacy of our methods on a number of examples, involving comparison with related methodologies in the literature. Current work concentrates towards extending our analysis to controlled systems, thus co-designing a controller and a certificate at the same time.
APPENDIX
1 Proof of Theorem 4
We aim at finding a bound on the discretisation gap $L(\theta,\xi)-L(\theta,\tilde{\xi})$ so that, for sufficiently small loss evaluated on the time-discretised approximations $L(\theta,\tilde{\xi})$, we also achieve a negative loss on the continuous trajectories $L(\theta,\xi)$.
$L(\theta,\xi)-L(\theta,\tilde{\xi})=l^{\Delta}(\theta,\xi)-l^{\Delta}(\theta,\tilde{\xi})$
$=\max_{x\in\xi}\evaluated{\frac{dB}{dt}}_{x}-\max_{k=1,\dots,M}\frac{B(x(t_{k}))-B(x(t_{k-1}))}{t_{k}-t_{k-1}}.$ (19)
Replacing the first maximisation with one between time instances, and exchanging the order of the $\max$ operators, (19) is equal to
$\max_{k=1,\dots,M}\max_{t\in[t_{k-1},t_{k}]}\evaluated{\frac{dB}{dt}}_{x(t)}$ (20)
$\qquad-\max_{k=1,\dots,M}\frac{B(x(t_{k}))-B(x(t_{k-1}))}{t_{k}-t_{k-1}},$
$\leq\max_{k=1,\dots,M}\Bigg{[}\max_{t\in[t_{k-1},t_{k}]}\evaluated{\frac{dB}{dt}}_{x(t)}-\frac{B(x(t_{k}))-B(x(t_{k-1}))}{t_{k}-t_{k-1}}\Bigg{]}.$ (21)
We can now replace the difference term with an integral, so that (21) becomes equivalent to
$\max_{k=1,\dots,M}\frac{\int_{t_{k-1}}^{t_{k}}\max_{t\in[t_{k-1},t_{k}]}\evaluated{\frac{dB}{dt}}_{x(t)}-\evaluated{\frac{dB}{dt}}_{x(\tau)}~{}\mathrm{d}\tau}{t_{k}-t_{k-1}}.$
Letting $\mathfrak{L}=\mathcal{M}_{B}\mathcal{L}_{f}+\mathcal{M}_{f}\mathcal{L}_{B}$ (refer to (14) for the definition of the various constants), the previous derivations lead to
$L(\theta,\xi)-L(\theta,\tilde{\xi})$
$\leq\max_{k=1,\dots,M}\frac{\int^{t_{k}}_{t_{k-1}}\|x(\tau)-\max_{t\in[t_{k-1},t_{k}]}x\|\mathfrak{L}~{}\mathrm{d}\tau}{t_{k}-t_{k-1}}$
$\leq\max_{k=1,\dots,M}\mathfrak{L}\frac{\int_{t_{k}}^{t_{k-1}}\mathcal{M}_{f}(t_{k}-t_{k-1})~{}\mathrm{d}\tau}{t_{k}-t_{k-1}},$ (22)
$=\max_{k=1,\dots,M}\mathfrak{L}\int_{t_{k}}^{t_{k-1}}\mathcal{M}_{f}~{}\mathrm{d}\tau=\overline{t}\mathfrak{L}\mathcal{M}_{f}.$ (23)
where the second inequality is since $\sup_{x}\|f(x)\|\leq\mathcal{M}_{f}$, and the last one since $\overline{t}=\max_{k=1,\dots,M}(t_{k}-t_{k-1})$.
This results then to a discretisation gap as in (14). By Theorem 3, and noticing that violating the conditions with $\psi^{\Delta}_{d}$ in place of $\psi^{\Delta}$, is equivalent to $L(\theta^{\star},\tilde{\xi})>-d$, we have
$\mathbb{P}^{N}\left\{\{\tilde{\xi}^{i}\}_{i=1}^{N}\colon\mathbb{P}\{\tilde{\xi}\colon L(\theta^{\star},\tilde{\xi})>-d\}\leq\varepsilon(C_{N},\beta,N)\right\}\geq 1-\beta.$
Since $L(\theta^{\star},\xi)\leq L(\theta^{\star},\tilde{\xi})+d$, this then implies that
$\mathbb{P}^{N}\left\{\{\tilde{\xi}^{i}\}_{i=1}^{N}\colon\mathbb{P}\{\xi\colon L(\theta^{\star},\xi)>0\}\leq\varepsilon(C_{N},\beta,N)\right\}\geq 1-\beta,$
thus concluding the proof. ∎

### v2

@@
 In addition, we provide probabilistic generalisation guarantees from discrete samples of continuous trajectories, to unseen continuous ones.
-Numerical investigations demonstrate the efficacy of our approach and contrast it with related results in the literature.
+Numerical results demonstrate the efficacy of our approach and contrast it with related results in the literature.
 For the purpose of Open Access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
-1 Introduction
+I Introduction
 Ensuring the safety of continuous-time dynamical systems is of critical importance in an increasingly autonomous world [1, 2, 3].
-As it is often infeasible to model system behaviour precisely, and making direct use of system data to verify behaviour is of interest [1, 12].
-A technique to verify properties of dynamical systems involves discretising the state space [3], under approximation guarantees, and verifying the resulting model.
-Alternatively, the use of certificates [2, 18] allows one to analyse directly the continuous-state system.
-These certificates map system’s states to real values, and exhibit certain properties that are relevant for analysis: here in particular we construct safety certificates for continuous-time systems, but extensions to more complex certificates are possible  [19].
+It is often infeasible to model system behaviour precisely, thus making direct use of system data to verify behaviour is of interest [4, 5].
+A technique for verifying properties of dynamical systems involves discretising the state space [6], under approximation guarantees, and verifying the resulting model.
+Alternatively, the use of certificates [7, 8] allows one to directly analyse the continuous-state system.
+These certificates map the system’s states to real values, and exhibit certain properties that are relevant for analysis: here, in particular, we construct safety certificates for continuous-time systems, but extensions to the more complex certificates in [3] are possible.
 There are a number of techniques for synthesising such certificates.
 In the case that an exact model is known, one can use a polynomial function as a certificate to formulate a convex sum-of-squares problem [9].
-Recent work in this area investigated the use of neural networks as certificates [8], which represent a class of general function approximators.
-When obtaining an exact model is infeasible, we turn to data-driven techniques.
+Recent work in this area investigated the use of neural networks (which represent a class of general function approximators) as certificates [1].
+Obtaining a model of the system, however, is generally a difficult task.
+It requires domain-specific knowledge and, since parts of the system may often not be known exactly, such a model will only constitute an approximation.
+To alleviate these issues, we turn to model-free data-driven techniques.
 One method for employing data in certificate synthesis is through the use of state pairs (i.e.
@@
 Such techniques are investigated in [2] for deterministic systems, and in [10] for stochastic systems.
-Both these works make use of the techniques in [11, 14] to bound the distance between what is referred to as a robust program, and its sample based counterpart.
-As discussed in [14, Remark 3.9], such techniques exhibit an exponential growth in the dimension of the sampling space (here the state space), and also require access to the underlying probability distribution.
-Alternatively, one can consider using entire trajectories as samples, hence only requiring access to realistic runs of the system.
-This technique is explored in [19] for discrete-time systems.
-This work leverages the so-called scenario approach [5, 10] and the notion of compression [13, 7], and provides a constructive algorithm for the general pick-to-learn framework [16], to provide a probably approximately correct (PAC) bound on the correctness of certificates for newly sampled trajectories.
-These techniques are restricted to discrete-time systems, since they perform calculations on entire trajectories, which is infeasible in continuous-time.
-Hence, here we develop techniques for continuous-time systems by analysing samples from complete trajectories and bounding the difference between the safety of these approximations and the complete trajectories.
-Our main contribution is the synthesis of neural barrier certificates for continuous-time dynamical systems, accompanied by PAC generalisation guarantees.
-These certificates are useful per se as a proxy for safety, and open the road for their use in control synthesis, an area of ongoing research.
+Both these works make use of the techniques in [11, 12] to bound the distance between what is referred to as a robust program, and its sample-based counterpart.
+As discussed in [12, Remark 3.9], such techniques exhibit an exponential growth in the dimension of the sampling space (here the state space), and also require access to the underlying probability distribution.
+Alternatively, one can consider using entire trajectories as samples, hence only requiring access to runs compatible with the dynamics of the system.
+This is explored in [3] for discrete-time systems.
+This work leverages the so-called scenario approach [13, 14] and the notion of compression [15, 16], and provides a constructive algorithm for the general pick-to-learn framework [17], to provide a probably approximately correct (PAC) bound on the correctness of certificates for newly sampled discrete trajectories.
+However, these guarantees are no longer valid when it comes to continuous-time systems.
+Here, we extend these developments to allow constructing a safety/barrier certificate based on a finite number of discretised trajectories of the continuous-time system; however, we provide guarantees on the safety of a new unseen continuous-time trajectory.
+The latter does not follow from the results in [3].
+Our contributions can be summarised as follows:
+  • We develop novel theoretical results that extend the results of [3] (which was limited in scope to discrete-time systems) to verify entire continuous trajectories, using discretised approximations for computations;
+  • We complement the results in [2], following an alternative approach for data-driven certificate synthesis which, however, does not scale exponentially with respect to the dimension of the underlying state space;
+  • Our approach is entirely data-driven, avoiding the use of satisfiability modulo theories (SMT) solvers, which are computationally expensive and require a system model.
 Notation.
 We use $\{\xi_{k}\}_{k=0}^{K}$ to denote a sequence indexed by $k\in\{0,1,\dots,K\}$.
-$B\models\psi$ defines condition satisfaction i.e., it evaluates to true if the quantity $B$ on the left satisfies the condition $\psi$ on the right, e.g., $x=1\models x>0$ evaluates to true and $x=-1\models x>0$ evaluates to false.
-Using ⊧$\not\models$̸ represents the logical inverse of this (i.e., condition dissatisfaction).
+$B\models\psi$ defines condition satisfaction, i.e., it evaluates to true if the quantity $B$ on the left satisfies the condition $\psi$ on the right.
+Using $\not\models$ represents the logical inverse of this (i.e., condition dissatisfaction).
 By $(\forall\xi\in\Xi)B\models\psi(\xi)$ we mean that some quantity $B$ satisfies a condition $\psi$ which, in turn, depends on some parameter $\xi$, for all $\xi\in\Xi$.
-We use $\xi_{[0,k]}$ to refer to a (possibly infinite) subsequence $\{\xi_{0},\dots,\xi_{k}\}$ of a sequence.
-2 Certificates
-In this work, we focus on barrier certificates, but emphasise that our techniques naturally extend to the more complex certificates in [19].
-To this end, we begin by defining a dynamical system, before considering the barrier certificate and associated safety property it verifies.
-2.1 Continuous-Time Dynamical Systems
-We consider a bounded state space $X\subseteq\mathbb{R}^{n}$, and a dynamical system whose evolution starts at an initial state $x(0)\in X_{I}$, where $X_{I}\subseteq X$ denotes the set of possible initial conditions.
-From an initial state, we uncover a finite trajectory, i.e., a continuous sequence of states $\xi=\{x(t)\}_{t\in[0,T]}$, where $T\in\mathbb{R}$ and $x(t)\colon t\rightarrow\mathbb{R}^{n}$, by following the dynamics
+We use $\xi_{[0,k]}$ to refer to a subsequence $\{\xi_{0},\dots,\xi_{k}\}$ of a sequence.
+II Safety and Barrier Certificates
+We focus on barrier certificates as the tool to verify safety; however, our techniques naturally extend to more complex certificates, as in [3].
+Consider a bounded state space $X\subseteq\mathbb{R}^{n}$, and a continuous-time dynamical system whose evolution starts at an initial state $x(0)\in X_{I}$, where $X_{I}\subseteq X$ denotes the set of possible initial conditions.
+From an initial state, we unravel a finite trajectory, i.e., a continuous sequence of states $\xi=\{x(t)\}_{t\in[0,T]}$, where $T\in\mathbb{R}$ and $x(t)\colon[0,T]\rightarrow\mathbb{R}^{n}$, by following the dynamics
 $\dot{x}(t)=f(x(t)).$ (1)
-We only require $f\colon X\rightarrow\mathbb{R}^{n}$ to be Lipschitz continuous, thus enforcing unique solutions.
+We only require $f\colon X\rightarrow\mathbb{R}^{n}$ to be Lipschitz continuous, thus allowing for existence and uniqueness of solutions.
 The set of all possible trajectories $\Xi\subseteq X_{I}\times X^{(0,T]}$ is then the set of all trajectories starting from the initial set $X_{I}$.
-In Section 3, we discuss how to use a finite set of trajectories to build safety certificates, and accompany them with generalization guarantees with respect to their validity for any trajectory.
-To allow for the synthesis of such certificate, as discussed in Section 4, we must store and perform calculations on such trajectories that play the role of samples.
-For this reason, we consider a finite $T$, and take samples along the continuous trajectory.
-Hence, we define the time-discretised trajectory
+In Section III, we discuss how to use a finite set of trajectories to build safety certificates, and accompany them with generalisation guarantees with respect to their validity for a new, unseen trajectory.
+We will also consider time-discretised versions of continuous-time trajectories.
+To this end, define the time-discretised trajectory
 $\tilde{\xi}=\{x(t)\}_{t\in\{0,t_{1},\dots,t_{M}\}}\in\tilde{\Xi}\subseteq X_{I}\times X^{M},$ (2)
 for $M$ sampled time steps $t_{1},\dots,t_{M}$.
-2.2 Safety Property
-We use $\phi(\xi)$ to refer to the safety property under study, evaluated on a trajectory $\xi\in\Xi$, defined as follows.
 Property 1 (Safety)
-Consider (1), and let $X_{I},X_{U}\subset X$ with $X_{I}\cap X_{U}=\varnothing$ denote an initial and an unsafe set, respectively. If, for all $\xi\in\Xi$,
-$\phi(\xi)\vcentcolon=\forall t\in[0,T],x(t)\notin X_{U},$
-holds, then we say that $\phi$ encodes a safety property. $\Xi$ denotes the set of trajectories consistent with (1) and with initial state contained within $X_{I}$.
-By the definition of $\phi$, it follows that verifying a system exhibits the safety property is equivalent to verifying all trajectories emanating from the initial set avoid the unsafe set for all time instances, until horizon $T$.
-2.3 Barrier Certificate
-We now define the relevant criteria necessary for a certificate $B$ to verify a safety property.
-We fix a time horizon $T<\infty$ and assume that $B$ is continuous, so that when considering the supremum/infimum of $B$ over $X$ (already assumed to be bounded) or over some of its subsets, this is well-defined.
+Consider (1), and let $X_{I},X_{U}\subset X$ with $X_{I}\cap X_{U}=\varnothing$ denote an initial and an unsafe set, respectively.
+We say that $\phi$ encodes a safety property for a trajectory $\xi\in\Xi$ if,
+$\phi(\xi)\vcentcolon=\forall t\in[0,T],x(t)\notin X_{U}.$
+By the definition of $\phi$, it follows that verifying that a trajectory exhibits the safety property is equivalent to verifying that a trajectory emanating from the initial set avoids the unsafe set for all time instances, until the horizon $T<\infty$.
+We now define the relevant criteria for a certificate $B$ to verify a safety property.
+Assume that $B$ is continuous, so that when considering the supremum/infimum of $B$ over $X$ (already assumed to be bounded) or over some of its subsets, this is well-defined.
 Consider:
@@
 $B(x)>0,\forall x\in X_{U},$ (4)
-$\evaluated{\frac{dB}{dt}}_{x\in\xi}<\frac{1}{T}\Big{(}\inf_{x\in X_{U}}B(x)-\sup_{x\in X_{I}}B(x)\Big{)},$ (5)
-where we may expand the Lie derivative as
-$\frac{dB}{dt}=\frac{\partial B}{\partial x}\frac{dx}{dt}=\frac{\partial B}{\partial x}f(x),$ (6)
-and hence recognise that this depends on the system dynamics $f(x)$. We place the following assumption on this Lie derivative, necessary for our analysis later.
-Assumption 1 (Lie Derivative is Lipschitz)
+[…]
+where $\frac{dB}{dt}=\frac{\partial B}{\partial x}f(x)$, and hence recognise that this depends on the system dynamics $f(x)$.
+Assumption 1 (Lipschitz Derivative)
 Assume that $\frac{dB}{dt}$ is Lipschitz continuous.
-Notice that even if $\inf_{x\in X_{U}}B(x)-\sup_{x\in X_{I}}B(x)>0$, i.e., in the case where the last condition encodes an increase of $B$ along the system trajectories, the system still avoids entering the unsafe set.
+Note that even if $\inf_{x\in X_{U}}B(x)-\sup_{x\in X_{I}}B(x)>0$, i.e., if the last condition encodes an increase of $B$ along the system trajectories, the system still avoids entering the unsafe set.
 This is established in the proof of Proposition 1 below.
-A graphical representation of these conditions is found in Figure 3.
-The level set (with dashed line) is the set obtained when the certificate value is 00.
-The decrease condition then means that we never leave the sublevel set.
 Denote by $\psi^{s}$ the conjunction of (3) and (4), and by $\psi^{\Delta}(\xi)$ the property in (5).
@@
 $B\models\psi^{s}\wedge(\forall\xi\in\Xi)B\models\psi^{\Delta}(\xi).$ (6)
-Proof 2.1.
+Proof:
 It suffices to show that satisfaction of (5) implies safety. Integrating (5) up to $t\leq T$, we obtain
@@
 The latter implies that all trajectories that start in $X_{I}$ avoid entering the unsafe set $X_{U}$, thus concluding the proof.
-To synthesise one of these deterministic certificates, we require complete knowledge of the behaviour $f$ of the dynamical system, to allow us to evaluate the Lie derivative $\frac{dB}{dt}$.
+∎
+To synthesise a certificate, we require complete knowledge of the behaviour $f$ of the dynamical system, to allow us to evaluate the derivative $\frac{dB}{dt}$.
 This may be impractical, and we therefore use data-driven techniques to learn a certificate.
-3 Data-Driven Certificates
+III Data-Driven Certificate Synthesis
 For our analysis, we will treat the initial state as random, distributed according to $\mathbb{P}$ (an appropriate probability space is defined; we gloss the technical details here in the interest of space).
@@
 the initial set $X_{I}$).
-To obtain our sample set, we consider $N$ independent and identically distributed (i.i.d.) initial conditions, sampled according to probability distribution $\mathbb{P}$, namely $\{x^{i}(0)\}_{i=1}^{N}\sim\mathbb{P}^{N}$.
-Initializing the dynamics from each of these initial states, we unravel a set of continuous-time trajectories $\{\xi^{i}\}_{i=1}^{N}$.
+We consider $N$ independent and identically distributed (i.i.d.) initial conditions, sampled according to $\mathbb{P}$, namely $\{x^{i}(0)\}_{i=1}^{N}\sim\mathbb{P}^{N}$.
+Initializing the dynamics from each of these initial states, we unravel a set of continuous-time trajectories $\{\xi^{i}\}_{i=1}^{N}$, also referred to as a multi-sample.
 Since there is no stochasticity in the dynamics, we can equivalently say that trajectories (generated from the random initial conditions) are distributed according to the same probabilistic law; hence, with a slight abuse of notation, we write $\xi\sim\mathbb{P}$.
-We impose the following assumption.
-Assumption 1 (Non-concentrated Mass).
+We impose the following assumption, standard in the scenario approach [16].
+Intuitively, this assumption rules out degenerate problem instances, where we could select the same sample multiple times with nonzero probability.
+Assumption 2 (Non-concentrated Mass)
 Assume that $\mathbb{P}\{\xi\}=0$, for any $\xi\in\Xi$.
-3.1 Problem Statement
-Since we are now dealing with a sample-based problem, we will construct probabilistic certificates and hence probabilistic guarantees on the satisfaction of a given property.
-Denote by $B_{N}$ a barrier certificate, we introduce the subscript $N$ to emphasize that this certificate is constructed on the basis of time-discretised sampled trajectories $\{\tilde{\xi}^{i}\}_{i=1}^{N}$.
-Problem 2 (Probabilistic Property Guarantee).
-Consider $N$ sampled trajectories, and fix a confidence level $\beta\in(0,1)$. We seek a property violation level $\epsilon\in(0,1)$ such that
-$\mathbb{P}^{N}$$\big{\{}\{\tilde{\xi}^{i}\}_{i=1}^{N}\in\tilde{\Xi}^{N}:~{}$
-$\mathbb{P}\{\xi\in\Xi\colon B_{N}\not\models\psi^{s}\wedge\psi^{\Delta}(\xi)\}\leq\varepsilon\big{\}}\geq 1-\beta.$ (10)
-In words, finding a solution to Problem 2 requires determining an $\epsilon\in(0,1)$, such that with confidence at least $1-\beta$, the probability that $B_{N}$ does not satisfy the condition $\psi^{s}\wedge\psi^{\Delta}(\xi)$ for another sampled continuous-time trajectory $\xi\in\Xi$ is at most equal to that $\epsilon$.
-As such, with a certain confidence, a certificate $B_{N}$ trained on the basis of $N$ sampled trajectories, will remain a valid certificate with probability at least $1-\epsilon$.
-Note that the outer probability in (10) (as well as in similar statements below) refers to the selection of discretized trajectories, $\{\tilde{\xi}^{i}\}_{i=1}^{N}$ (as these are used for training), while the inner probability refers to the selection of a continuous time one $\xi$, as this captures the desired generalization properties.
-3.2 Probabilistic Guarantees
+Since we are now dealing with a sample-based problem, we construct probabilistic certificates and hence probabilistic guarantees on the satisfaction of a given property.
+III-A The Discrete-Time Case
+Designing safety certificates that enjoy probabilistic guarantees when it comes to new unseen trajectories has been recently considered in [3], however, for discrete-time systems.
+We review the main developments in this direction, extending these to continuous-time ones.
 Consider a mapping $\mathcal{A}$ such that $B_{N}=\mathcal{A}(\{\xi^{i}\}_{i=1}^{N})$ as an algorithm that, based on $N$ samples, returns a certificate $B_{N}$.
-We call as compression set of such an algorithm any subset of the input that returns the same certificate.
-That is, a sample-subset $\mathcal{C}_{N}\subseteq\{\xi^{i}\}_{i=1}^{N}$ is a compression set if $\mathcal{A}(\mathcal{C}_{N})=\mathcal{A}(\{\xi^{i}\}_{i=1}^{N})$.
+We call a compression set of such an algorithm any subset of the input multi-sample that returns the same certificate.
+That is, a sample subset $\mathcal{C}_{N}\subseteq\{\xi^{i}\}_{i=1}^{N}$ is a compression set if $\mathcal{A}(\mathcal{C}_{N})=\mathcal{A}(\{\xi^{i}\}_{i=1}^{N})$.
 In Algorithm 1, we provide a specific synthesis procedure through which $\mathcal{A}$ (and hence the certificate $B_{N}$) can be constructed.
-This algorithm is adapted from [19], with appropriate modifications to allow for continuous-time dynamics; we discuss this in the next section.
-Using the results of [19], and if we have access to samples of the time-derivative $f(x)$, we can provide a guarantee over the time-discretised trajectories, as stated next.
-Theorem 3 (Probabilistic Guarantees [19]).
+This algorithm is extended from [3] to continuous-time systems and is discussed in the next section.
+Theorem 1 (Probabilistic Guarantees [3])
 Consider Assumption 2, and let $B_{N}$ and $\mathcal{C}_{N}$ be the certificate and compression set, respectively, returned by Algorithm 1.
-Fix $\beta\in(0,1)$, and for $k<N$, let let $\varepsilon(k,\beta,N)$ be the (unique) solution to the polynomial equation in the interval $[k/N,1]$
+Fix $\beta\in(0,1)$, and for $k<N$, let $\varepsilon(k,\beta,N)$ be the (unique) solution to the polynomial equation in the interval $[k/N,1]$
 $\frac{\beta}{2N}\sum_{m=k}^{N-1}$ $\frac{\binom{m}{k}}{\binom{N}{k}}(1-\varepsilon)^{m-N}$
@@
 where $C_{N}=|\mathcal{C}_{N}|$ is the cardinality of the compression set.
-Unfortunately, this does not provide us with a guarantee on the property satisfaction for the continuous-time trajectories, i.e.
-our trajectory may violate the safety property between sampled states $x(t_{i})$ and $x(t_{i+1})$, as the inner probability refers to a choice of $\tilde{\xi}$ as opposed to $\xi$.
-If the Lipschitz continuity in Assumption 1 holds, we can additionally offer guarantees on newly sampled continuous-time trajectories, based only on time-discretised trajectories, and approximate derivatives.
-In this case, we require a tightening of the derivative condition in (5) by some value $d\in\mathbb{R}$, to ensure that the decrease condition is maintained between sample times.
-Denote this condition (over discretized trajectories) as $\psi^{\Delta}_{d}(\tilde{\xi})$, defined by the inequality
+Remark 1
+Theorem 1 is not applicable to continuous-time trajectories.
+In particular, the inner probability refers to the probability of a new time-discretised trajectory $\tilde{\xi}$ being (un)safe.
+As a result, states along the state trajectory $\xi$ of the continuous-time system under study may violate the safety property (chiefly, between sampled states $x(t_{i})$ and $x(t_{i+1})$).
+III-B Extension to the Continuous-Time Case
+[…]
 $\max_{k=1,\dots,M}$ $\frac{B(x(t_{k}))-B(x(t_{k-1}))}{t_{k}-t_{k-1}}$ (11)
-$<\frac{1}{T}\Big{(}\inf_{x\in X_{U}}B(x)-\sup_{x\in X_{I}}B(x)\Big{)}-d.$
-Define by $L_{B}$ and $L_{f}$ the Lipschitz constants of the certificate derivative $\frac{\partial B_{\theta^{\star}}}{\partial x}$ and of the dynamics $f(x)$ respectively, and by $\mathcal{M}_{B},\mathcal{M}_{f}$ bounds on their norms, namely $\sup_{x}\|\frac{\partial B_{\theta^{\star}}}{\partial x}\|\leq\mathcal{M}_{B}$ and $\sup_{x}\|f(x)\|\leq\mathcal{M}_{f}$.
-Then, the value $d$ is defined as
-$d=\overline{t}\mathcal{M}_{f}\left(\mathcal{M}_{B}\mathcal{L}_{f}+\mathcal{M}_{f}\mathcal{L}_{B}\right),$ (14)
-where $\overline{t}=\max_{k=1,\dots,M}(t_{k}-t_{k-1})$.
-Theorem 4 (Continuous-Time Guarantees).
+$<\frac{1}{T}\Big{(}\inf_{x\in X_{U}}B(x)-\sup_{x\in X_{I}}B(x)\Big{)}-d,$
+where $d\in\mathbb{R}$ is a tightening parameter.
+Define by $\mathcal{L}_{B}$ and $\mathcal{L}_{f}$ the Lipschitz constants of the certificate derivative $\frac{\partial B_{\theta^{\star}}}{\partial x}$ and of the dynamics $f(x)$ respectively, by $\mathcal{M}_{B},\mathcal{M}_{f}$ bounds on their norms, i.e.
+$\sup_{x}\|\frac{\partial B_{\theta^{\star}}}{\partial x}\|\leq\mathcal{M}_{B}$ and $\sup_{x}\|f(x)\|\leq\mathcal{M}_{f}$, and for $\overline{t}=\max_{k=1,\dots,M}(t_{k}-t_{k-1})$, define $d$ as
+$d=\overline{t}\mathcal{M}_{f}\left(\mathcal{M}_{B}\mathcal{L}_{f}+\mathcal{M}_{f}\mathcal{L}_{B}\right).$ (12)
+Theorem 2 (Continuous-Time Guarantees)
 Consider the conditions of Theorem 1, Assumption 1 and (12). Then,
@@
 $\mathbb{P}\{\xi\in\Xi\colon B_{N}\not\models\psi^{s}\wedge\psi^{\Delta}(\xi))\}\leq\varepsilon(C_{N},\beta,N)\big{\}}\geq 1-\beta.$
-The proof of this is achieved by bounding the difference between the safety of the continuous-time trajectories and their time-discretised approximations, and is in the Appendix.
-4 Certificate Synthesis
-In order to learn a barrier certificate from samples, we consider a neural network, a well-studied class of function approximators that generalize well to a given task.
-Denote all tunable neural network parameters by a vector $\theta$.
+The proof of this is achieved by bounding the difference between the safety of the continuous-time trajectories and their time-discretised approximations; see Appendix.
+IV Certificate Synthesis
+To learn a barrier certificate from samples, we consider a neural network, a well-studied class of function approximators that generalise well to a given task, although any other function approximation mechanism could be chosen instead.
+We consider a network with $N_{\mathrm{layers}}$ layers, and $N_{\mathrm{neurons}}$ neurons per layer.
+Denote all tunable neural network parameters by a vector $\theta$ $\in\mathbb{R}^{N_{\mathrm{layers}}N_{\mathrm{neurons}}}$.
 We then have that our certificate $B_{N}$ depends on $\theta$.
-For the results of this section, we simply write $B_{\theta}$ and drop the dependency on $N$ to ease notation.
+For the results of this section, we write $B_{\theta}$ and drop the dependency on $N$ to ease notation.
 Algorithm 1 Certificate Synthesis and Compression Set Computation
 1:function $\mathcal{A}(\theta,\mathcal{D})$
-2:     Set $k\leftarrow 1$triangleright Initialise iteration index
-3:     Set $\mathcal{C}\leftarrow\varnothing$triangleright Initialise compression set
-4:     Fix $L_{1}<L_{0}$ with $|L_{1}-L_{0}|>\eta$triangleright $\eta$ is any fixed tolerance
+2:  Set $k\leftarrow 1$ $\triangleright$ Initialise iteration index
+3:  Set $\mathcal{C}\leftarrow\varnothing$ $\triangleright$ Initialise compression set
+4:  Fix $L_{1}<L_{0}$ with $|L_{1}-L_{0}|>\eta$ $\triangleright$ $\eta$ is any fixed tolerance
 5:  while $l^{s}(\theta)>0$ do $\triangleright$ While sample-independent state loss is non-zero
-6:         $g\leftarrow\nabla_{\theta}l^{s}(\theta)$triangleright Gradient of loss function
-7:         $\theta\leftarrow\theta-\alpha g$triangleright Step in the direction of sample-independent gradient
+6:   $g\leftarrow\nabla_{\theta}l^{s}(\theta)$ $\triangleright$ Gradient of loss function
+7:   $\theta\leftarrow\theta-\alpha g$ $\triangleright$ Step in the direction of sample-independent gradient
 8:
 9:  while $|L_{k}-L_{k-1}|>\eta$ do $\triangleright$ Iterate until tolerance is met
-10:         $\mathcal{M}\leftarrow\{\tilde{\xi}\in\mathcal{D}\colon L(\theta,\tilde{\xi})\geq\max_{\tilde{\xi}\in\mathcal{C}}L(\theta,\tilde{\xi})\}$triangleright Find maximal samples with loss greater than compression set loss
-11:         $\overline{g}_{\mathcal{M}}\leftarrow\{\nabla_{\theta}L(\theta,\tilde{\xi})\}_{\tilde{\xi}\in\mathcal{M}}$triangleright Subgradients of loss function for $\tilde{\xi}\in\mathcal{M}$
-12:         $\overline{\xi}_{\mathcal{C}}\in\operatorname*{arg\,max}_{\tilde{\xi}\in\mathcal{C}}L(\theta,\tilde{\xi})$triangleright Find a sample with maximum loss from $\mathcal{C}$
-13:         $\overline{g}_{\mathcal{C}}\leftarrow\nabla_{\theta}L(\theta,\overline{\xi}_{\mathcal{C}})$triangleright Approximate subgradient of loss function for $\tilde{\xi}=\overline{\xi}_{\mathcal{C}}$
+[…]
+11:   $\overline{g}_{\mathcal{M}}\leftarrow\{\nabla_{\theta}L(\theta,\tilde{\xi})\}_{\tilde{\xi}\in\mathcal{M}} \triangleright$ Subgradients of loss function for $\tilde{\xi}\in\mathcal{M}$
+[…]
+13:   $\overline{g}_{\mathcal{C}}\leftarrow\nabla_{\theta}L(\theta,\overline{\xi}_{\mathcal{C}})$ $\triangleright$ Approximate subgradient of loss function for $\tilde{\xi}=\overline{\xi}_{\mathcal{C}}$
 14:
 15:   if $\exists\overline{g}\in\overline{g}_{\mathcal{M}}\colon\langle\overline{g},\overline{g}_{\mathcal{C}}\rangle\leq 0\wedge\overline{g}\neq 0$  then $\triangleright$ If there is a misaligned subgradient (take the maximum if multiple)
-16:              $\theta\leftarrow\theta-\alpha\overline{g}$triangleright Step in the direction of misaligned subgradient
-17:              $\mathcal{C}\leftarrow\mathcal{C}\cup\{\overline{\xi}\}$triangleright Update compression set with sample corresponding to $\overline{g}$
+16:     $\theta\leftarrow\theta-\alpha\overline{g}$ $\triangleright$ Step in the direction of misaligned subgradient
+17:     $\mathcal{C}\leftarrow\mathcal{C}\cup\{\overline{\xi}\}$ $\triangleright$ Update compression set with sample corresponding to $\overline{g}$
 18:   else
-19:              $\theta\leftarrow\theta-\alpha\overline{g}_{\mathcal{C}}$triangleright Step in the direction of approximate subgradient
+19:     $\theta\leftarrow\theta-\alpha\overline{g}_{\mathcal{C}}$ $\triangleright$ Step in the direction of approximate subgradient
 20:
-21:         $L_{k}\leftarrow\min\left\{L_{k-1},\max_{\tilde{\xi}\in\mathcal{D}}L(\theta,\tilde{\xi})\right\}$triangleright Update “running” loss value
-22:         $k\leftarrow k+1$triangleright Update iteration index
+[…]
+22:   $k\leftarrow k+1$ $\triangleright$ Update iteration index
 23:  return $\theta,\mathcal{C}$
-Figure 1: Graphical Representation of Algorithm 1.
-4.1 Certificate and Compression Set Computation
-We seek to minimize a loss function that encodes the barrier certificate conditions, with respect to the neural network parameters.
-To this end, for a $\tilde{\xi}\in\tilde{\Xi}$ and parameter vector $\theta$, let
+IV-A Certificate Synthesis Algorithm
+We seek to minimise a loss function that encodes the barrier certificate conditions, with respect to the neural network parameters. For $\tilde{\xi}\in\tilde{\Xi}$ and parameter vector $\theta$, let
 $L(\theta,\tilde{\xi})=l^{\Delta}(\theta,\tilde{\xi})+l^{s}(\theta),$ (14)
 represent an associated loss function comprised of sample-dependent loss $l^{\Delta}$, and sample-independent loss $l^{s}$
-$l^{s}(V_{\theta})$ :$\vcentcolon=\frac{1}{|\mathcal{X}_{I}|}\sum_{x\in\mathcal{X}_{I}}\max\{0,V_{\theta}(x)\}$
-$+\frac{1}{|\mathcal{X}_{U}|}\sum_{x\in\mathcal{X}_{U}}\max\{0,-V_{\theta}(x)\}.$
-$l^{\Delta}(\theta,\tilde{\xi})$ :$\vcentcolon=-\frac{1}{T}\Big{(}\inf_{x\in\mathcal{X}_{U}}B(x)-\sup_{x\in\mathcal{X}_{I}}B(x)\Big{)}$
-$+\max_{k=1,\dots,M}\frac{B(x(t_{k}))-B(x(t_{k-1}))}{t_{k}-t_{k-1}}.$
-To instantiate these functions we consider a discrete set of grid-points on each sub-domain: $\mathcal{X}_{\overline{I}}$ is the set of points in the domain but outside the goal region, and $\mathcal{X}_{U}$ is the set of points in the unsafe set.
+$l^{s}(\theta)$ $\vcentcolon=\frac{1}{|\mathcal{X}_{I}|}\sum_{x\in\mathcal{X}_{I}}\max\{0,B_{\theta}(x)\}$
+$+\frac{1}{|\mathcal{X}_{U}|}\sum_{x\in\mathcal{X}_{U}}\max\{0,\delta-B_{\theta}(x)\}.$
+[…]
+$+\max_{k=1,\dots,M}\frac{B_{\theta}(x(t_{k}))-B_{\theta}(x(t_{k-1}))}{t_{k}-t_{k-1}},$
+To instantiate these functions we consider a discrete set of $N_{\mathrm{states}}$ grid-points on each sub-domain: $\mathcal{X}_{I}$ is the set of points in the initial set, and $\mathcal{X}_{U}$ is the set of points in the unsafe set.
 These points are generated densely enough across the domain of interest, and hence offer an accurate approximation.
-Since these samples do not require access to the dynamics, we consider them separate to the sample-set $\{\tilde{\xi}^{i}\}_{i=1}^{N}$.
-Consider the first summation in the sample-independent loss, if $B(x)<0$ then $\max\{0,B_{\theta}(x)\}=0$, i.e., no loss is incurred, implying satisfaction of (3).
-Under a similar reasoning, the other integral accounts for (4), respectively.
-We impose the next mild assumption, a sufficient condition for termination of our algorithm.
-Assumption 5 (Minimizers’ Existence).
-For any $\{\tilde{\xi}\}_{i=1}^{N}$, and any non-empty $\mathcal{D}\subseteq\{\tilde{\xi}\}_{i=1}^{N}$, the set of minimizers of $\max_{\tilde{\xi}\in\mathcal{D}}L(\theta,\tilde{\xi}),$ is non-empty.
-Note that, for a sufficiently expressive neural network, we can find a certificate $B$ which satisfies the state constraints and hence has a sample-independent loss of zero.
-Algorithm 1 provides an inexact subgradient methodology to minimize the loss function, and to iteratively construct a compression set $\mathcal{C}$ (initially empty; see step $3$).
-We explain the main steps of this algorithm with reference to Figure 1.
-After an arbitrary parameter initialization, we follow the subgradient associated with the current sample in $\mathcal{C}$ (“blue” dot labeled by $3$).
-At this point, this step becomes inexact, as there would exist another sample resulting in a higher loss (“green” point labeled by $2$).
+Since they do not require access to the dynamics, we consider them separately from $\{\tilde{\xi}^{i}\}_{i=1}^{N}$.
+[…]
+The following mild assumption ensures a minimiser of the loss functions exists and hence our algorithm terminates.
+Assumption 3 (Minimisers’ Existence)
+[…]
+Algorithm 1 provides an inexact subgradient methodology to minimise the loss function, and to iteratively construct a compression set $\mathcal{C}$ (initially empty; see step $3$).
+We explain the main steps of this algorithm with reference to Figure 1, where each sample gives rise to a concave triangular constraint.
+After an arbitrary initialization, we follow the subgradient associated with the worst case sample and add it to the compression set $\mathcal{C}$ (step 14141515, point $M_{1}$).
+When iterates get to point $M_{2}$, the subgradient step becomes inexact, as for the same parameter there exists a sample resulting in a higher loss (see asterisk).
 Such a sample is in $\mathcal{M}$, step $9$ of Algorithm 1.
-However, the algorithm does not “jump” to the green point, as the condition in step $13$ of the algorithm is not yet satisfied.
-As such the algorithm performs inexact subgradient descent steps up to point $5$; this is the first instance where the condition in step $13$ is satisfied (there exists another constraint with opposite slope) and hence the algorithm “jumps” to a point with higher loss and subgradient of opposite sign.
+However, the algorithm does not “jump”, as the condition in step $13$ of the algorithm is not yet satisfied.
+Graphically, this is since the $M_{2}$ and the red asterisk are on a side of the respective constraint with the same slope.
+As such the algorithm performs inexact subgradient descent steps up to point $M_{3}$; this is the first instance where the condition in step $13$ is satisfied (there exists another constraint with opposite slope) and hence the algorithm “jumps” to a point with higher loss and subgradient of opposite sign.
+This procedure is then repeated as shown in the figure, with the red line indicating the iterates’ path.
 The “jumps” serve as an exploration step to investigate the non-convex landscape, while their number corresponds to the cardinality of the returned compression set.
-Such a procedure can be thought of as a constructive procedure for the general framework presented recently in [16] to construct compression sets.
-In some cases, the parameter returned by Algorithm 1 may result in a value of the loss function greater than $-d$, and hence mean we are unable to verify safety.
-To achieve a lower loss, we make use of a sample-and-discarding procedure [6, 20], and introduce Algorithm 2 as an outer loop around Algorithm 1.
-This procedure leads to a lower loss, however the samples that are discarded have to be added to the compression set.
-Theorems 3 and 4 then apply, but with the compression set returned by Algorithm 2 instead.
-To terminate our algorithm we require knowledge of $d$, but cannot calculate $d$ until we find $\theta^{\star}$. To resolve this, we propose two different approaches.
+This can be thought of as a constructive procedure for the general framework presented in [17] to construct compression sets.
+The computational complexity of one iteration of the main loop is $\mathcal{O}((M+N_{\mathrm{states}})\cdot N\cdot N_{\mathrm{layers}}\cdot N_{\mathrm{neurons}}^{3})$.
+We can upper bound the number of iterations by $\frac{L_{0}-L^{\star}}{\eta}$, where $L^{\star}$ is the value of the loss associated with one of the minimisers.
+To terminate our algorithm we require knowledge of $d$ (which depends on $\theta^{\star}$).
+To resolve this, we propose two approaches.
   1. 1.
-    At every iteration $j$ calculate $d_{j}$ using the current best parameters $\theta_{j}$, terminate when $\max_{i}\left[L(\theta_{j},\tilde{\xi}^{i})\right]<d_{j}$.
+    At every iteration $j$ calculate $d_{j}$ using the current best parameters $\theta_{j}$, terminate if $\max_{i}\left[L(\theta_{j},\tilde{\xi}^{i})\right]<-d_{j}$.
   2. 2.
-    Choose a parameter set $\Theta$, and take the supremum across the set to find an upper bound on $\mathcal{L}_{B},\mathcal{M}_{B}$, use these to calculate $d$.
-To determine Lipschitz constants for neural networks we refer the reader to [4, 9, 22].
+    Choose a parameter set, and consider the supremum across that set to determine an upper bound on $\mathcal{L}_{B},\mathcal{M}_{B}$. Use these upper bounds to calculate $d$.
+The second option is likely to be conservative but removes the need for calculating Lipschitz constants in the loop, hence is computationally more efficient.
+To determine Lipschitz constants for neural networks we refer to [18, 19].
+[…]
+Figure 1: Graphical illustration of Algorithm 1.
+IV-B Sample Discarding and Compression Computation
+Due to the tightening introduced, to minimise the loss function, we would like its value to be lower than or equal to d-d.
+However, in some cases, the parameter returned by Algorithm 1 may result in a loss value greater than d-d, thus preventing us from verifying safety.
+To ensure the loss is below that threshold, we employ a sampling-and-discarding procedure [20, 21], and introduce Algorithm 2 as an outer loop around Algorithm 1.
+Algorithm 2 iterates as long the desired loss value is not met (step $4$), and progressively discards samples (step $7$).
+The samples that are discarded are the ones identified as compression by Algorithm 1 (step $6$), as removing these is bound to improve the solution.
 Algorithm 2 Compression Set Update with Discarding
@@
 2:Set $\widetilde{\mathcal{C}}\leftarrow\varnothing\triangleright$ Initialise compression set
-3:Set $\mathcal{D}\leftarrow\{\tilde{\xi}^{i}\}_{i=1}^{N}$triangleright Initialise “running” samples
+3:Set $\mathcal{D}\leftarrow\{\tilde{\xi}^{i}\}_{i=1}^{N}$ $\triangleright$ Initialise “running” samples
 4:while $(\max_{\tilde{\xi}\in\mathcal{D}}l^{\Delta}(\theta,\tilde{\xi})>-d)\bigvee(l^{s}(\theta)>0)$ do
-5:     $\theta,\mathcal{C}\leftarrow$ $\mathcal{A}(\theta,\mathcal{D})$triangleright Call Algorithm 1
-6:     $\widetilde{\mathcal{C}}\leftarrow\widetilde{\mathcal{C}}\cup\mathcal{C}$triangleright Update $\widetilde{\mathcal{C}}$
-7:     $\mathcal{D}\leftarrow\mathcal{D}\setminus\widetilde{\mathcal{C}}$triangleright Discard $\widetilde{\mathcal{C}}$ from $\mathcal{D}$
+5:  $\theta,\mathcal{C}\leftarrow$ $\mathcal{A}(\theta,\mathcal{D})$ $\triangleright$ Call Algorithm 1
+6:  $\widetilde{\mathcal{C}}\leftarrow\widetilde{\mathcal{C}}\cup\mathcal{C}$ $\triangleright$ Update $\widetilde{\mathcal{C}}$
+7:  $\mathcal{D}\leftarrow\mathcal{D}\setminus\widetilde{\mathcal{C}} \triangleright$ Discard $\widetilde{\mathcal{C}}$ from $\mathcal{D}$
 8:return $\theta$, $\widetilde{\mathcal{C}}$
-5 Numerical Results
-We consider constructing a safety certificate for the nonlinear, two-dimensional jet engine model as considered in [15],
-$\dot{x}_{1}(t)=-x_{2}(t)-\frac{3}{2}x_{1}^{2}(t)-\frac{1}{2}x_{1}^{3}(t),\;\;\dot{x}_{2}(t)=x_{1}(t).$ (17)
-In Figure 3 we provide a graphical representation of the dynamics, sub-domains under study, the 00-level set produced by our certificate, and the level sets calculated by the methods in [15] (one lower bounding the unsafe set, the other upper bounding the initial set).
-We used $5$ independent repetitions (each with different multi-samples) of $1,000$ sampled trajectories and $367$ seconds of computation time (standard deviation $139$s), to obtain $\varepsilon=0.01492$ (standard deviation $0.00140$) with confidence to $0.99$.
+The samples that are discarded need to be added to the compression set (step $6$); intuitively this is the case as if we repeat the procedure using only the compression samples we need to follow the same solution path.
+A direct byproduct of Algorithms 1 and 2 is that they calculate a compression set, whose cardinality $C_{N}$ is in turn used in Theorem 2 to provide the desired probabilistic guarantees.
+The cardinality of that set corresponds to the number of “jumps” in Algorithm 1 plus the number of discarded samples in Algorithm 2.
+This cardinality is not necessarily greater for higher dimensional problems, but is rather case dependent and depends on the complexity of the problem.
+For example, a problem where some trajectories approach or even enter the unsafe set presents a more challenging synthesis problem than one where trajectories all move in the opposite direction to the unsafe set, thus we expect the former to have a larger compression set even if the problem is smaller in dimension.
+V Numerical Results
+We consider constructing a safety certificate for the nonlinear, two-dimensional jet engine model as considered in [2], The codebase is available at https://github.com/lukearcus/fossil_scenario
+$\dot{x}_{1}(t)=-x_{2}(t)-\frac{3}{2}x_{1}^{2}(t)-\frac{1}{2}x_{1}^{3}(t),\;\;\dot{x}_{2}(t)=x_{1}(t),$ (15)
+using a time horizon $T=5$.
+Figure 3 provides a graphical representation of the dynamics, subdomains under study, the 0-level set produced by our certificate, and the level sets calculated by the methods in [2] (one lower bounding the unsafe set, the other upper bounding the initial set).
+We used $5$ independent repetitions (each with different multi-samples) of $1,000$ sampled trajectories and $367$ seconds of computation time (standard deviation $139$s), to obtain $\varepsilon=0.01492$ (standard deviation $0.00140$) with confidence $0.99$.
 The methodology of [2] required $257149$ state pair samples and $5123$ seconds (standard deviation $449$s) of computation time to compute a barrier certificate with the same confidence (however, this holds deterministically).
-We estimate the Lipschitz constants using the methods in [23].
+We estimate the Lipschitz constants using the methodology in [22], noting that convergence is guaranteed asymptotically.
 Figure 3 contains a 3D plot of the certificate.
-Figure 2: Phase plane plot, initial and unsafe set for (17).
-The zero-level set for our certificate is dashed; level sets that bound the initial and unsafe sets (i.e.
-$\gamma$- and $\lambda$- level sets) in [15] are dotted.
+Figure 2: Phase plane plot, initial and unsafe set for (15). The zero-level set for our certificate is dashed; level sets that bound the initial and unsafe sets in [2] are dotted.
 Figure 3: Surface plot of the safety/barrier certificate, generated by our techniques, for the system of Figure 3.
-Beyond these numerical results, we briefly discuss the theoretical differences between our approaches.
+Beyond these numerical results, we briefly discuss the theoretical differences between our approach and [2].
 The results in [2] offer a guarantee that, with a certain confidence, the safety property is always satisfied, in contrast to Theorem 2 where we provide such guarantees in probability (up to a quantifiable risk level $\varepsilon$).
-However, these “always” guarantees, albeit very useful, come with some challenges.
-Firstly, they are not applicable when part of the initial set is unsafe.
-Secondly, they implicitly require some knowledge of the underlying probability distribution.
-Finally, they are bound to an exponential growth with the dimension of the state space.
+However, these “always” guarantees, although very useful, come with some challenges.
+Firstly, they are not applicable when part of the initial set is unsafe, whereas we can still bound the probability of a new trajectory being safe even if some of the sampled trajectories were unsafe.
+Second, they implicitly require knowledge of the underlying probability distribution to instantiate their confidence bound and scale exponentially with respect to the state space dimension.
 Related to this last point, we performed a comparison on the following four-dimensional system taken from [1].
@@
 $\dot{x}_{2}(t)$ $=\cos(x_{4}(t)),$
-$\dot{x}_{3}(t)$ $=0.01\sqrt{|x_{1}(t)|},$
-$\dot{x}_{4}(t)$ $=-x_{1}(t)-x_{2}(t)^{2}+\sin(x_{4}(t)).$
-We calculate that the approach of [15] requires at least $10^{19}$ samples to return a confidence lower bounded by at most $10^{-30}$, which is not practically useful.
-In contrast, our techniques with only $100$ samples obtain a risk level of $\varepsilon=0.21450$ (standard deviation $0.00910$), confidence $1-10^{-5}$.
-6 Conclusion
+[…]
+$\dot{x}_{4}(t)$ $=-x_{1}(t)-x_{2}(t)^{2}+\sin(x_{4}(t)),$
+using $T=4$.
+We applied also the approach of [2] which, with $10^{19}$ samples, returned a confidence $10^{-30}$, an uninformative result as it is close to zero.
+With only $100$ samples our techniques obtain a risk level $\varepsilon=0.21450$ (standard deviation $0.00910$), confidence $1-10^{-5}$ (close to one).
+VI Conclusion
 We have proposed a method for synthesis of neural-network certificates for continuous-time dynamical systems, based only on a finite number of trajectories from a system.
 Our numerical experiments demonstrate the efficacy of our methods on a number of examples, involving comparison with related methodologies in the literature.
-Current work concentrates towards extending our analysis to controlled systems, thus co-designing a controller and a certificate at the same time.
+Current work concentrates towards extending to controlled systems, thus co-designing a controller and a certificate at the same time.
 APPENDIX
-1 Proof of Theorem 4
-We aim at finding a bound on the discretisation gap $L(\theta,\xi)-L(\theta,\tilde{\xi})$ so that, for sufficiently small loss evaluated on the time-discretised approximations $L(\theta,\tilde{\xi})$, we also achieve a negative loss on the continuous trajectories $L(\theta,\xi)$.
-$L(\theta,\xi)-L(\theta,\tilde{\xi})=l^{\Delta}(\theta,\xi)-l^{\Delta}(\theta,\tilde{\xi})$
-$=\max_{x\in\xi}\evaluated{\frac{dB}{dt}}_{x}-\max_{k=1,\dots,M}\frac{B(x(t_{k}))-B(x(t_{k-1}))}{t_{k}-t_{k-1}}.$ (19)
-Replacing the first maximisation with one between time instances, and exchanging the order of the $\max$ operators, (19) is equal to
-$\max_{k=1,\dots,M}\max_{t\in[t_{k-1},t_{k}]}\evaluated{\frac{dB}{dt}}_{x(t)}$ (20)
+I Proof of Theorem 2
+[…]
+[…]
+[…]
+Replace the first maximisation with one between time instances, and exchange the order of the $\max$ operators,
+[…]
 $\qquad-\max_{k=1,\dots,M}\frac{B(x(t_{k}))-B(x(t_{k-1}))}{t_{k}-t_{k-1}},$
-$\leq\max_{k=1,\dots,M}\Bigg{[}\max_{t\in[t_{k-1},t_{k}]}\evaluated{\frac{dB}{dt}}_{x(t)}-\frac{B(x(t_{k}))-B(x(t_{k-1}))}{t_{k}-t_{k-1}}\Bigg{]}.$ (21)
-We can now replace the difference term with an integral, so that (21) becomes equivalent to
-$\max_{k=1,\dots,M}\frac{\int_{t_{k-1}}^{t_{k}}\max_{t\in[t_{k-1},t_{k}]}\evaluated{\frac{dB}{dt}}_{x(t)}-\evaluated{\frac{dB}{dt}}_{x(\tau)}~{}\mathrm{d}\tau}{t_{k}-t_{k-1}}.$
+[…]
+We can now replace the difference term with an integral,
+[…]
 Letting $\mathfrak{L}=\mathcal{M}_{B}\mathcal{L}_{f}+\mathcal{M}_{f}\mathcal{L}_{B}$ (refer to (12) for the definition of the various constants), the previous derivations lead to
-$L(\theta,\xi)-L(\theta,\tilde{\xi})$
+[…]
 $\leq\max_{k=1,\dots,M}\frac{\int^{t_{k}}_{t_{k-1}}\|x(\tau)-\max_{t\in[t_{k-1},t_{k}]}x\|\mathfrak{L}~\mathrm{d}\tau}{t_{k}-t_{k-1}}$